OpenStudy (anonymous):

GIVE MEDALS! What is the equation of the line that passes through (4, -1) and (-2, 3)? 2x - 3y - 5 = 0 2x + 3y - 5 = 0 -2x + 3y - 5 = 0

OpenStudy (anonymous):

@jziggy

OpenStudy (jziggy):

change in y over change in x to get the slope so, (3 - (-1))/(-2 - 4) which simplifies to -2/3

OpenStudy (jziggy):

So you now have y = -2x/3 + b where we need to solve for b by plugging in one of the coordinates. -1 = -8/3 + b b = 5/3

OpenStudy (jziggy):

So, y = -2x/3 + 5/3 and now we just need to do a little gymnastics to make it look right

OpenStudy (jziggy):

Multiply both sides of the equation by 3 to get 3y = -2x + 5

OpenStudy (jziggy):

and then 3y + 2x - 5 = 0